OpenStudy (anonymous):

Functions r and s are defined as all real numbers x as follows r(x)=2x^2 and s(x)=3x. Write the expressions for(r+s)(x) and (r-s)(x) and evaluate (r x s) (2).

OpenStudy (anonymous):

Take the equations r(x) and s(x) and substitute them into the variables r and s in the new equations f(x) = (r(x)+s(x)) and g(x) = (r(x)-s(x)). Once you have those equations, evaluate both when x = 2.

OpenStudy (anonymous):

f(x) = r(x) + s(x) = 2x^2 + 3x, so f(2) = r(2) + s(2) = 2*(2)^2 + 3*(2) = 14